consider a Technician at KT Spares and workshop. The worker is required o disassemble the wheel from the wheel assembly, remove the old Tyre and store it permanently, replace it with the new Tyre after inspection and reassemble it to the wheel assembly and inspect. the technician spends minutes to remove wheel, 5 minutes to replace the Tyre, 4 minutes to fit the new Tyre and 6 Minutes to assemble the wheel. the company operates on one day shift per weekday except weekends for four weeks in a month. The Technician is paid a gross of $5000/ month at the current demand of one Tyre replacement every 15 minutes.

Requiresments:

a) Calculate the labor productivity in units per month
b) Determine the number of vehicles queued at the end of the day and month
c) Is this business profitable if the tyre replacement is priced at $60/tyre and $45/tyre on overheads
1
Expert's answer
2017-11-24T14:00:06-0500
The technician spends 0 minutes to remove wheel, 5 minutes to replace the Tyre, 4 minutes to fit the new Tyre and 6 Minutes to assemble the wheel. The company operates on one day shift per weekday except weekends for four weeks in a month. The Technician is paid a gross of $5000/ month at the current demand of one Tyre replacement every 15 minutes.
a) The labor productivity in units per month is:
60*8*20/15 = 640 units per worker per month.
b) Determine the number of vehicles queued at the end of the day is:
60*8/15/4 = 8 vehicles.
At the and of the month:
8*20 = 160 vehicles.
c) If the tyre replacement is priced at $60/tyre and $45/tyre on overheads, then the amount of profit is: TP = TR - TC = P*Q - (VC + FC) = 60*640 - 45*640 - 5000 = $4600.
So, this business is profitable.